# Learning Where to Learn: Training Distribution Selection for Provable OOD Performance

### v1

@@ Abstract
Out-of-distribution (OOD) generalization remains a fundamental challenge in machine learning. Models trained on one data distribution often experience substantial performance degradation when evaluated on shifted or unseen domains. To address this challenge, the present paper studies the design of training data distributions that maximize average-case OOD performance. First, a theoretical analysis establishes a family of generalization bounds that quantify how the choice of training distribution influences OOD error across a predefined family of target distributions. These insights motivate the introduction of two complementary algorithmic strategies: (i) directly formulating OOD risk minimization as a bilevel optimization problem over the space of probability measures and (ii) minimizing a theoretical upper bound on OOD error. Last, the paper evaluates the two approaches across a range of function approximation and operator learning examples. The proposed methods significantly improve OOD accuracy over standard empirical risk minimization with a fixed distribution. These results highlight the potential of distribution-aware training as a principled and practical framework for robust OOD generalization.
@@ 1 Introduction
Out-of-distribution (OOD) generalization is a core obstacle for machine learning algorithms, particularly when the test-time data distribution differs significantly from that of training. Standard methods, such as empirical risk minimization (ERM) which focuses solely on minimizing the average loss over the training distribution, often overlook the complexities of distribution shift. This oversight can lead to substantial performance degradation when models are exposed to unseen or shifted domains. Such covariate shifts are common in scientific machine learning tasks such as solving infinite-dimensional inverse problems and operator learning for partial differential equations (PDEs).
This work proposes a framework for optimizing the training distribution to improve average-case OOD performance. Rather than accepting the limitations imposed by a fixed training distribution, the paper systematically investigates how the choice of training data influences generalization to target distributions. Guided by both theoretical insights and practical considerations, we develop a principled and computationally feasible strategy for learning where to learn: selecting the training data distribution that ensures robust performance across a specified family of test distributions.
Motivation.
In nonlinear inverse problems like electrical impedance tomography (EIT), neural networks known as neural operators are used to learn maps between spaces of functions, e.g., from boundary voltage patterns to internal conductivity [19, 39, 42]. However, their success heavily depends on how the training data is chosen. In EIT, the input data comes from the so-called Dirichlet-to-Neumann (DtN) map $\Lambda_{a}$. It encodes how the voltage applied on the boundary of a domain leads to current flow and ultimately reveals information about the unknown conductivity $a$—the output data—inside the domain. We use input data pairs $\{(f_{n},g_{n})\}$ for training, where $f_{n}$ is a voltage pattern applied at the boundary (a Dirichlet condition), and $g_{n}=\Lambda_{a}f_{n}$ is the resulting current response (a Neumann condition). These pairs are sampled from a probability distribution $\nu$ over the space of input voltages that we can choose freely. Crucially, this distribution $\nu$ has a significant impact on how well the model generalizes to new, unseen patterns. In Fig. 1, we explore this phenomenon using the Neural Inverse Operator (NIO) [39] and its variant; see SM DE for details.
Figure 1: Two conductivity samples in EIT. The true conductivity is on the left of each panel, followed by predictions from models trained on four Dirichlet boundary condition distributions, $\nu_{1}$$\nu_{4}$. Top row: in-distribution predictions; bottom row: OOD predictions. Refer to SM DE for details.
@@
The primary goal of this paper is to understand and optimize the role of training data distributions in improving OOD generalization for regression problems. Operator learning is a driving motivation, with a focus on challenging inverse problems such as EIT and parametrized PDE solution operators such as Darcy flow. Our main contributions are now listed.
@@
    Lipschitz-based OOD bounds. We develop generalization bounds that connect OOD error to the Lipschitz continuity of the learned operator and quantitatively characterize how the Wasserstein mismatch between train and test distributions impact average-case performance.
@@
    Algorithmic strategies for training distribution optimization. Motivated by the theory, we propose two label-dependent algorithms for actively optimizing the training distribution: bilevel optimization targeting average-case OOD risk and a computationally efficient alternative that minimizes an upper risk bound; both outperform nonadaptive ERM.
@@
    Empirical demonstration of distributional effects in operator learning. Our numerical experiments demonstrate that the choice of training data distribution significantly influences generalization error in both forward and inverse operator learning tasks, consistent with the theoretical insights. Additionally, we showcase the effectiveness and efficiency of the two proposed algorithms in actively optimizing the training data distribution.
@@
OOD generalization [37] connects to active learning, optimal experimental design (OED), and robust domain generalization. Active learning [16, 49, 55] improves model performance by selectively querying informative samples based on criteria like uncertainty or diversity. However, these methods operate reactively on unlabeled datasets during training, without addressing the proactive optimization of the initial training distribution. In contrast, our work emphasizes the systematic design of training distributions to improve robustness to distribution shifts, particularly in operator learning tasks like inverse problems and surrogate modeling. OED similarly focuses on distribution selection but within a statistical or Bayesian framework [6, 20, 21, 26]. Classical methods, such as A-optimality and D-optimality, aim to reduce estimation error, while Bayesian and sequential designs adaptively balance exploration and exploitation. Although these approaches focus on parameter inference or uncertainty reduction, our work departs by targeting training distribution optimization for minimizing OOD errors, which is not a commonly considered design principle in OED. Finally, our work aligns with robust domain generalization efforts, including distributionally robust optimization (DRO) [33, 45] and invariant risk minimization (IRM) [5, 28, 34, 56]. While DRO minimizes worst-case risk over predefined groups and IRM enforces invariance across domains, both approaches focus on robustness during training rather than optimizing the training data distribution itself. In contrast, we propose a novel framework that actively designs training distributions, guided by theoretical bounds and bilevel optimization, to address OOD challenges in regression problems.
@@
This paper is organized as follows. Sec. 2 provides the necessary background. Sec. 3 establishes motivating theoretical results. Sec. 4 proposes two implementable data distribution selection algorithms. Sec. 5 performs numerical studies and Sec. 6 gives concluding remarks. Details about derivations and numerical implementation are deferred to the supplementary material (SM).
@@ 2 Preliminaries
This section sets the notation, provides essential background on key mathematical concepts, and reviews the framework of operator learning. SM A contains additional reference material.
@@
This paper works with an abstract input space $(\mathcal{U},\langle\cdot,\cdot\rangle_{\mathcal{U}},\lVert{\,\cdot\,}\rVert_{\mathcal{U}})$ and output space $(\mathcal{Y},\langle\cdot,\cdot\rangle_{\mathcal{Y}},\lVert{\,\cdot\,}\rVert_{\mathcal{Y}})$, both assumed to be real separable Hilbert spaces. We often drop the subscripts on norms and inner products when the meaning is clear from the context. Unless otherwise stated, $\mathcal{U}$ or $\mathcal{Y}$ could be infinite-dimensional. We denote the Lipschitz constant of a map $F\colon\mathcal{U}\to\mathcal{Y}$ by $\operatorname{Lip}(F)\coloneqq\sup_{u\neq u^{\prime}}\lVert F(u)-F(u^{\prime})\rVert_{\mathcal{Y}}/\lVert u-u^{\prime}\rVert_{\mathcal{U}}$. For any element $\zeta$ of a metric space, the Dirac probability measure assigning unit mass to $\zeta$ is written as $\delta_{\zeta}$. For $p\in[1,\infty]$, (Bochner) $L^{p}_{\mu}(\mathcal{U};\mathcal{Y})$ spaces are defined in the usual way. Let $\mathsf{m}_{p}\colon\mathscr{P}(\mathcal{U})\to\mathbb{R}_{\geq 0}\cup\{\infty\}$ denote the uncentered $p$-th moment defined by $\mu\mapsto\mathsf{m}_{p}(\mu)\coloneqq\operatorname{\mathbb{E}}_{u\sim\mu}\lVert u\rVert_{\mathcal{U}}^{p}$. We say that a probability measure $\mu$ belongs to $\mathscr{P}_{p}(\mathcal{U})$ if $\mathsf{m}_{p}(\mu)<\infty$. We often work with Gaussian measures $\mathcal{N}(m,\mathcal{C})\in\mathscr{P}_{2}(\mathcal{U})$, where $m\in\mathcal{U}$ is the mean and $\mathcal{C}\colon\mathcal{U}\to\mathcal{U}$ is the symmetric trace-class positive semidefinite covariance operator; see [14, App. A.3] for details in the infinite-dimensional setting.
@@
Consider probability measures $\mu$ and $\nu$ on a metric space $(X,\mathsf{d})$ with finite $p$-th moments, where $p\in[1,\infty)$. Define the set of all couplings $\Gamma(\mu,\nu)\coloneqq{\mathopen{}\mathclose{{}\left\{\gamma\in\mathscr{P}(X\times X)\,\middle|\,(\pi^{1})_{\#}\gamma=\mu,\;(\pi^{2})_{\#}\gamma=\nu}\right\}}$, where $\pi^{1}(x,y)=x$, $\pi^{2}(x,y)=y$, and $\mathscr{P}(X\times X)$ is the set of probability measures over $X\times X$. The $p$-Wasserstein distance between $\mu$ and $\nu$ is $\mathsf{W}_{p}(\mu,\nu)\coloneqq(\inf_{\gamma\in\Gamma(\mu,\nu)}\int_{X\times X}\mathsf{d}(x,y)^{p}\,\gamma(dx,dy))^{1/p}$.
@@
We will use the notion of random measure [27] to quantify OOD generalization error in Secs. 34.
@@
Operator learning provides a powerful framework for approximating mappings $\mathcal{G}^{\star}\colon\mathcal{U}\to\mathcal{Y}$ between function spaces. It plays an increasingly pivotal role for accelerating forward and inverse problems in the computational sciences [9, 29, 30, 41]. As in traditional regression, the learned operator $\mathcal{G}_{\widehat{\theta}}$ is parametrized over a set $\Theta$ and minimizes an empirical risk
@@
The Hilbert norm in (2.1) and discretization-invariant architectures (e.g., DeepONet [36], Fourier Neural Operator (FNO) [32], NIO [39]) are points of departure from classical function approximation. In (2.1), $\{u_{n}\}\sim\nu^{\otimes N}$, where $\nu$ is the training distribution over input functions. The choice and properties of $\nu$ are crucial. An inadequately selected $\nu$ may result in $\mathcal{G}_{\widehat{\theta}}(u)$ being a poor approximation to $\mathcal{G}^{\star}(u)$ for $u\sim\nu^{\prime}$, where $\nu^{\prime}\neq\nu$ [1, 2, 10, 15, 31, 40, 43, 52]. The present work aims to improve such OOD generalization in both function approximation and operator learning settings.
3 Lipschitz theory for out-of-distribution generalization
This section studies how loss functions behave under distribution shifts and the implications of such shifts on average-case OOD accuracy. The assumed Lipschitz continuity of the target map and approximation class play a central role in the analysis.
@@ 3 Lipschitz theory for out-of-distribution generalization
Proposition 3.1 (OOD generalization).
@@
See SM B for the proof. Provided that the maps $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are Lipschitz, Prop. 3.1 shows that the functional $\mu\mapsto\operatorname{\mathbb{E}}_{u\sim\mu}\lVert\mathcal{G}_{1}(u)-\mathcal{G}_{2}(u)\rVert$ is globally Lipschitz continuous in the $\mathsf{W}_{1}$ metric. Similarly, the functional $\mu\mapsto\operatorname{\mathbb{E}}_{u\sim\mu}\lVert\mathcal{G}_{1}(u)-\mathcal{G}_{2}(u)\rVert^{2}$ is locally Lipschitz continuous in the $\mathsf{W}_{2}$ metric; in particular, it is Lipschitz on sets with uniformly bounded second moments. We remark that the Lipschitz map hypotheses in Prop. 3.1 are invoked for simplicity. Hölder continuity is also compatible with Wasserstein distances, and the preceding result can be adapted to hold under such regularity.
@@
OOD generalization performance may be quantified in several ways. This paper chooses average-case accuracy with respect to the law $\mathbb{Q}\in\mathscr{P}(\mathscr{P}(\mathcal{U}))$ of a random probability measure $\nu^{\prime}\sim\mathbb{Q}$. For a ground truth map $\mathcal{G}^{\star}$ and any hypothesis $\mathcal{G}$, we strive for
@@
to be small. Other approaches to quantify OOD accuracy are possible, such as worst-case analysis.
@@
We now instantiate Cor. 3.2 in the concrete context of Gaussian mixtures.
Example 3.3 (Gaussian mixtures).
Consider the random Gaussian measure $\nu^{\prime}=\mathcal{N}(m,\mathcal{C})\sim\mathbb{Q}\in\mathscr{P}_{2}(\mathscr{P}_{2}(\mathcal{U}))$, where $(m,\mathcal{C})\sim\Pi$ and $\Pi$ is the mixing distribution over the mean (with marginal $\Pi_{1}$) and covariance (with marginal $\Pi_{2}$). Define the Gaussian mixture $\overline{\nu}\coloneqq\operatorname{\mathbb{E}}_{(m,\mathcal{C})\sim\Pi}\mathcal{N}(m,\mathcal{C})\in\mathscr{P}_{2}(\mathcal{U})$ via duality. Let $\overline{\mu}\coloneqq\operatorname{\mathbb{E}}_{(m^{\prime},\mathcal{C}^{\prime})\sim\Pi^{\prime}}\mathcal{N}(m^{\prime},\mathcal{C}^{\prime})$ be another Gaussian mixture with mixing distribution $\Pi^{\prime}$ (with marginals denoted by $\Pi_{1}^{\prime}$ and $\Pi_{2}^{\prime}$). By linearity of expectation and Fubini’s theorem, $\mathscr{E}_{\mathbb{Q}}(\mathcal{G})=\mathscr{E}_{\delta_{\overline{\nu}}}(\mathcal{G})$; that is, average-case accuracy with respect to $\mathbb{Q}$ is equivalent to accuracy with respect to the single test distribution $\overline{\nu}$. This fact and Cor. 3.2 yield $\mathscr{E}_{\mathbb{Q}}(\mathcal{G})\leq\operatorname{\mathbb{E}}_{u\sim\overline{\mu}}\lVert\mathcal{G}^{\star}(u)-\mathcal{G}(u)\rVert^{2}+c(\mathcal{G}^{\star},\mathcal{G},\overline{\mu},\overline{\nu})\mathsf{W}_{2}(\overline{\mu},\overline{\nu})$. The factor $c(\mathcal{G}^{\star},\mathcal{G},\overline{\mu},\overline{\nu})$ can be further simplified using the closed-form expression for second moments of Gaussian measures. For the $\mathsf{W}_{2}$ factor, Lem. B.1 in SM B and coupling arguments imply that
$\mathscr{E}_{\mathbb{Q}}(\mathcal{G})\leq\operatorname{\mathbb{E}}_{u\sim\overline{\mu}}\big{\lVert}\mathcal{G}^{\star}(u)-\mathcal{G}(u)\big{\rVert}_{\mathcal{Y}}^{2}+c\bigl{(}\mathcal{G}^{\star},\mathcal{G},\overline{\mu},\overline{\nu}\bigr{)}\sqrt{\mathsf{W}_{2}^{2}(\Pi_{1},\Pi_{1}^{\prime})+\mathsf{W}_{2}^{2}(\Pi_{2},\Pi_{2}^{\prime})}\,.$ (3.5)
The two $\mathsf{W}_{2}$ distances between the marginals may be further bounded with (A.1) in SM A under additional structural assumptions on the mixing distributions; see Rmk. B.2. We often invoke this example in numerical studies with $\mathbb{Q}=\sum_{k=1}^{K}w_{k}\delta_{\nu_{k}^{\prime}}$ (where $K\in\mathbb{N}$, i.e., a finite mixture), weights $\sum_{k=1}^{K}w_{k}=1$, Gaussian test measures $\{\nu_{k}^{\prime}\}$, and degenerate training mixture $\overline{\mu}\coloneqq\mathcal{N}(m_{0},\mathcal{C}_{0})$.

### v3

# Learning where to learn: Training data distribution optimization for scientific machine learning

@@ Abstract
In scientific machine learning, models are routinely deployed with parameter values or boundary conditions far from those used in training. This paper studies the learning-where-to-learn problem of designing a training data distribution that minimizes average prediction error across a family of deployment regimes. A theoretical analysis shows how the training distribution shapes deployment accuracy. This motivates two adaptive algorithms based on bilevel or alternating optimization in the space of probability measures. Discretized implementations using parametric distribution classes or nonparametric particle-based gradient flows deliver optimized training distributions that outperform nonadaptive designs. Once trained, the resulting models exhibit improved sample complexity and robustness to distribution shift. This framework unlocks the potential of principled data acquisition for learning functions and solution operators of partial differential equations.
@@ 1 Introduction
In scientific settings, machine learning models encounter data at deployment time that can differ substantially from the data seen during training. Standard empirical risk minimization (ERM), which minimizes an average loss over the fixed training distribution, may yield models that perform well in-sample yet degrade under distribution shift. This challenge is acute in operator learning for partial differential equations (PDEs) and infinite-dimensional inverse problems. For example, in data-driven electrical impedance tomography (EIT), neural operators are trained to map electrical patterns on the boundary to internal conductivities [32, 67, 75]. The input is a collection of voltage and corresponding current pairs that represent Dirichlet and Neumann boundary conditions, respectively. The probability distribution $\nu$ that boundary voltages are sampled from is a user choice. Crucially, the choice of $\nu$ shapes how well the trained model generalizes to new boundary measurement distributions. Fig. 1 illustrates this effect with a variant of the Neural Inverse Operator (NIO) architecture [67]; see SM EF for details.
These observations motivate the central premise of the present work: learning where to learn by designing an optimal training distribution $\nu$ for a prescribed family of deployment regimes. Guided by theoretical and practical considerations, this paper develops a principled and computationally feasible framework for optimizing $\nu$ in function regression and operator learning tasks. It establishes intelligent data acquisition as a crucial component of scientific machine learning (SciML) workflows.
Figure 1: Two conductivity samples in EIT. The true conductivity is on the left of each panel, followed by predictions from NIO models trained on four Dirichlet boundary condition distributions $\{\nu_{i}\}_{i=1}^{4}$. Top row: in-distribution predictions; bottom row: out-of-distribution predictions. See SM EF for details.
@@
Our main contributions are as follows.
@@
    Lipschitz-based distribution shift bounds. We develop quantitative bounds that connect out-of-distribution (OOD) error to the training error, Lipschitz continuity of the model and target map, and the Wasserstein mismatch between training and test distributions.
@@
    Algorithms for training distribution optimization. Motivated by the theory, we develop two adaptive algorithms that actively optimize the training distribution for general model classes: ($1$) a bilevel procedure that directly minimizes OOD error and ($2$) a computationally efficient alternating scheme that minimizes an upper bound on the OOD error. We implement ($1$) and ($2$) over the space of probability measures by considering (a) parametric training distribution families (e.g., Gaussian processes) and (b) nonparametric particle representations of the training distribution (e.g., that are updated with Wasserstein gradient flows).
@@
    Empirical performance in SciML tasks. We perform numerical experiments involving function approximation as well as forward and inverse operator learning for a range of elliptic, parabolic, and hyperbolic PDEs (EIT, Darcy flow, viscous Burgers’, and radiative transport). Consistent with our theory, the results show that the training data distribution strongly governs OOD deployment error. In the experiments, the two proposed adaptive algorithms efficiently optimize the training distribution, greatly reduce OOD error, and often outperform adaptive and nonadaptive baselines.
@@
The careful selection of sampling points for regression appears frequently in numerical analysis [11, 55]. Weighted least-squares regression with Christoffel sampling is one prominent example [21, 3]. One of the main findings of these works is that training on samples from the target test distribution is a suboptimal strategy in general (cf. Canatar et al. [16]). Instead, a change of measure is derived by minimizing an upper bound on average error. These ideas have their origin in numerical integration and Monte Carlo methods [15]. Some other related design strategies include greedy Bayesian herding [46] and kernel quadrature [31]. However, quadrature-based sampling does not necessarily deliver optimal performance for regression.
In SciML for PDEs, several adaptive sampling and reweighting strategies have emerged, particularly for unsupervised methods such as physics-informed neural networks (PINNs) [99]. Common approaches either reweight a fixed pool of training data samples [17, 64] or sample new data points from a density proportional to the PDE residual [37, 91]. Similar methods based on importance sampling, reweighting, or coreset selection have been applied to neural operators [97, 92]. Others use residual information to move samples directly [74] or iteratively expand training datasets [62]. The current work builds on these advances by operating at the level of probability distributions. This enables a wide range of parametrizations and discretizations beyond samples only and provides extra freedom to explore the design space.
For a discussion of broader connections to other areas of machine learning and statistics, see SM A.
@@
This paper is organized as follows. Sec. 2 provides the necessary background. Sec. 3 establishes motivating theoretical results (C1). Sec. 4 proposes two implementable data distribution selection algorithms (C2). Sec. 5 performs numerical studies (C3) and Sec. 6 gives concluding remarks. Proof and implementation details are deferred to the supplementary material (SM).
@@ 2 Preliminaries
This section sets the notation, provides essential background on key mathematical concepts, and reviews the framework of operator learning. SM B contains additional reference material.
@@
This paper works with an abstract input space $(\mathcal{U},\langle\cdot,\cdot\rangle_{\mathcal{U}},\lVert{\,\cdot\,}\rVert_{\mathcal{U}})$ and output space $(\mathcal{Y},\langle\cdot,\cdot\rangle_{\mathcal{Y}},\lVert{\,\cdot\,}\rVert_{\mathcal{Y}})$, both assumed to be real separable Hilbert spaces. We often drop the subscripts on norms and inner products when the meaning is clear from the context. Unless otherwise stated, $\mathcal{U}$ or $\mathcal{Y}$ could be infinite-dimensional. We denote the Lipschitz constant of a map $F\colon\mathcal{U}\to\mathcal{Y}$ by $\operatorname{Lip}(F)\coloneqq\sup_{u\neq u^{\prime}}\lVert F(u)-F(u^{\prime})\rVert_{\mathcal{Y}}/\lVert u-u^{\prime}\rVert_{\mathcal{U}}$. For any element $\zeta$ of a metric space, the Dirac probability measure assigning unit mass to $\zeta$ is written as $\delta_{\zeta}$. For $p\in[1,\infty]$, (Bochner) $L^{p}_{\mu}(\mathcal{U};\mathcal{Y})$ spaces are defined in the usual way. We say that a probability measure $\mu\in\mathscr{P}(\mathcal{U})$ belongs to $\mathscr{P}_{p}(\mathcal{U})$ if $\mathsf{m}_{p}(\mu)<\infty$, where $\mathsf{m}_{p}\colon\mathscr{P}(\mathcal{U})\to\mathbb{R}_{\geq 0}\cup\{\infty\}$ denotes the uncentered $p$-th moment defined by $\mu\mapsto\mathsf{m}_{p}(\mu)\coloneqq\operatorname{\mathbb{E}}_{u\sim\mu}\lVert u\rVert_{\mathcal{U}}^{p}$. We often work with Gaussian measures $\mathcal{N}(m,\mathcal{C})\in\mathscr{P}_{2}(\mathcal{U})$, where $m\in\mathcal{U}$ is the mean and $\mathcal{C}\colon\mathcal{U}\to\mathcal{U}$ is the symmetric trace-class positive semidefinite covariance operator; see Dashti and Stuart [23, App. A.3] for details in the infinite-dimensional setting.
@@
Consider probability measures $\mu$ and $\nu$ on a metric space $(X,\mathsf{d})$ with finite $p$-th moments, where $p\in[1,\infty)$. Define the set of all couplings $\Gamma(\mu,\nu)\coloneqq{\mathopen{}\mathclose{{\left\{\gamma\in\mathscr{P}(X\times X)\,\middle|\,\pi^{1}_{\#}\gamma=\mu,\;\pi^{2}_{\#}\gamma=\nu}}\right\}}$, where $\pi^{1}(x,y)=x$ and $\pi^{2}(x,y)=y$. The $p$-Wasserstein distance between $\mu$ and $\nu$ is $\mathsf{W}_{p}(\mu,\nu)\coloneqq(\inf_{\gamma\in\Gamma(\mu,\nu)}\int_{X\times X}\mathsf{d}(x,y)^{p}\,\gamma(dx,dy))^{1/p}$.
@@
We will use the notion of random measure [50] to quantify OOD error in Secs. 34.
@@
Operator learning provides a powerful framework for approximating mappings $\mathcal{G}^{\star}\colon\mathcal{U}\to\mathcal{Y}$ between function spaces. It plays an increasingly pivotal role for accelerating forward and inverse problems in the computational sciences [13, 52, 54, 68, 72, 90]. As in traditional regression, the learned operator $\mathcal{G}_{\widehat{\theta}}$ is parametrized over a set $\Theta$ and minimizes an empirical risk
@@
The Hilbert norm in (2.1) and discretization-invariant architectures (e.g., DeepONet [61], Fourier Neural Operator (FNO) [57], NIO [67]) are points of departure from classical function approximation. In (2.1), $\{u_{n}\}\sim\nu^{\otimes N}$, where $\nu$ is the training distribution over input functions. The choice and properties of $\nu$ are crucial. An inadequately selected $\nu$ may result in $\mathcal{G}_{\widehat{\theta}}(u)$ being a poor approximation to $\mathcal{G}^{\star}(u)$ for $u\sim\nu^{\prime}$, where $\nu^{\prime}\neq\nu$ [3, 14, 24, 56, 71, 76, 89]. The present work aims to improve accuracy under distribution shift in both function approximation and operator learning settings.
3 Lipschitz theory for out-of-distribution error bounds
This section studies how loss functions behave under distribution shifts and the implications of such shifts on average-case OOD accuracy. The assumed Lipschitz continuity of the target map and the elements in the approximation class play a central role in the analysis. Although our theoretical contributions build upon established tools, we are not aware of any existing work that presents these results in the precise form developed here.
@@ 3 Lipschitz theory for out-of-distribution error bounds
Proposition 3.1 (distribution shift error).
@@
See SM C for the proof. Provided that the maps $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are Lipschitz, Prop. 3.1 shows that the functional $\mu\mapsto\operatorname{\mathbb{E}}_{u\sim\mu}\lVert\mathcal{G}_{1}(u)-\mathcal{G}_{2}(u)\rVert$ is globally Lipschitz continuous in the $\mathsf{W}_{1}$ metric. Similarly, the functional $\mu\mapsto\operatorname{\mathbb{E}}_{u\sim\mu}\lVert\mathcal{G}_{1}(u)-\mathcal{G}_{2}(u)\rVert^{2}$ is locally Lipschitz continuous in the $\mathsf{W}_{2}$ metric; in particular, it is Lipschitz on sets with uniformly bounded second moments. We remark that the Lipschitz map hypotheses in Prop. 3.1 are invoked for simplicity. Hölder continuity is also compatible with Wasserstein distances, and the preceding result can be adapted to hold under such regularity.
@@
OOD performance may be quantified in several ways. This paper chooses average-case accuracy with respect to the law $\mathbb{Q}\in\mathscr{P}(\mathscr{P}(\mathcal{U}))$ of a random probability measure $\nu^{\prime}\sim\mathbb{Q}$. For a ground truth map $\mathcal{G}^{\star}$ and any hypothesis $\mathcal{G}$, we strive for
@@
to be small. An equivalent viewpoint in (D.2) from SM D.1 rewrites (3.3) as a single expectation against a mixture. Other approaches to quantify OOD error are possible, such as worst-case analysis. We choose average-case accuracy because it is more natural for learning from finite samples and because the linearity of (3.3) with respect to $\mathbb{Q}$ is useful for algorithm development.
@@
We instantiate Cor. 3.2 in the concrete context of Gaussian mixtures in SM C, Ex. C.1.